What is the total mass of these 70 cobalt atoms (in grams)? Express your answer to 2 significant figures.

TIP: To report an answer in scientific notation, enter it using the format "2.3E4", which means "2.3 x 104" (without the quotation marks)
1
Expert's answer
2019-09-23T03:45:42-0400

nCo=NCo/NA = 11.62E-23

mCo = n*M = 11.62E-23*59= 685.82E-23